# Multiple Stochastic Prompt Tuning for Practical Cross-Domain Few Shot Learning

### v1

@@ 4.2 MIST: Multiple Stochastic Prompt Tuning
Since CLIP is pretrained on web-scale data encompassing mainly natural images [24], the image encoder struggles to learn robust classwise features when faced with extreme domain and semantic shifts in the target dataset. In addition, intraclass diversity and interclass similarity in the images due to presence of all classes together, results in disjoint clusters in visual features from the same class. Consider an illustrative example of the EuroSAT dataset containing satellite images of various terrains. Here, distinct classes like “Highway” and “River” can have similar visual representatives in the few-shot training data, at the same time featuring diverse visuals from the same classes as shown in Figure 2. Existing prompt tuning approaches represent each class with single learnable prompts, implicitly assuming that classwise visual features form single clusters. However, such a strategy is insufficient to represent the disjoint visual clusters, which may result from such challenging settings. To illustrate this, we consider two representative prompt-tuning methods, MaPLe [15] and TCP [29] and show their t-SNE visualizations in Figure 2. We observe that the unimodal assumption is violated and the image embeddings of each class form multiple modes in the representation space.
To address this, inspired from [2, 1], we introduce a multiple prompt learning approach, where we represent each class with multiple prompt vectors. However, incorporating too many learnable prompts for each class is not desirable since: (i) Each class contains few training samples, and many text embeddings per class could lead to overfitting on individual datapoints, resulting in loss of class level representations, and (ii) it introduces additional learnable parameters and thus more computational overhead. As a balanced approach, we represent each class with two prompt vectors.
Formally, let the embeddings corresponding to the two prompts for a particular class $CLS_{k}$ be denoted as $\tilde{z}_{t}^{i}$, where, $i$ = 1, 2. Here, $\tilde{z}_{t}^{i}=\mathcal{F}_{t}(\tilde{X}_{t}^{i})$, where, $\tilde{X}_{t}^{i}=\{t_{SOS},\theta_{t}^{i},t_{1},...,t_{CLS_{k}},t_{EOS}\}$ is the $i^{th}$ text prompt for the class $CLS_{k}$. $\theta_{t}^{i}=\{\theta_{t1}^{i},\theta_{t2}^{i},...,\theta_{tm}^{i}\}$ denotes the $i^{th}$ set of learnable prompt vectors for that particular class.
@@
CoOp (IJCV’22) [32] $51.87$ $22.77$ $24.73$ 22.83 $30.55$
TaskRes (CVPR’23) [31] $64.67$ $19.70$ $36.57$ $10.97$ $32.98$
MaPLe (CVPR’23) [15] $73.30$ $27.50$ $51.53$ $14.60$ $41.73$
PromptSRC (ICCV’23) [16] $73.23$ $21.97$ 55.03 $14.37$ $41.15$
CLAP (CVPR’24) [25] $61.46$ $26.61$ $47.22$ $15.94$ $37.81$
TCP (CVPR’24) [29] $64.30$ $27.80$ $49.37$ $14.93$ $39.10$
@@
2-shots
CoOp (IJCV’22) [32] $66.00$ $21.87$ $37.97$ $14.43$ $35.07$
TaskRes (CVPR’23) [31] $68.83$ $23.13$ $39.27$ $10.83$ $35.52$
MaPLe (CVPR’23) [15] $78.07$ $31.90$ $67.17$ $16.27$ $48.35$
PromptSRC (ICCV’23) [16] $79.53$ $29.47$ $68.07$ $12.70$ $47.44$
CLAP (CVPR’24) [25] $70.63$ $34.79$ $60.13$ 16.43 $45.50$
TCP (CVPR’24) [29] $70.37$ 36.87 $62.63$ $15.63$ $46.38$
@@
4-shots
CoOp (IJCV’22) [32] $66.53$ $25.00$ $42.67$ $17.93$ $38.03$
TaskRes (CVPR’23) [31] $72.40$ $21.40$ $39.35$ $10.27$ $35.86$
MaPLe (CVPR’23) [15] $84.03$ $37.17$ $77.07$ 19.73 $54.50$
PromptSRC (ICCV’23) [16] $85.23$ $37.63$ $78.70$ $15.17$ $54.18$
CLAP (CVPR’24) [25] $76.43$ $34.37$ $65.11$ $18.98$ $48.72$
TCP (CVPR’24) [29] $76.77$ $37.37$ $67.97$ $17.07$ $49.80$
@@
8-shots
CoOp (IJCV’22) [32] $76.53$ $38.27$ $60.50$ $14.60$ $47.48$
TaskRes (CVPR’23) [31] $74.63$ $34.30$ $57.77$ $12.57$ $44.82$
MaPLe (CVPR’23) [15] $86.80$ $46.53$ $84.47$ $14.17$ $57.99$
PromptSRC (ICCV’23) [16] $88.37$ $42.47$ $86.80$ $14.93$ $58.14$
CLAP (CVPR’24) [25] $76.85$ $42.81$ $74.62$ $14.97$ $52.31$
TCP (CVPR’24) [29] $79.03$ $46.57$ $76.33$ $14.97$ $54.23$
@@
16-shots
CoOp (IJCV’22) [32] $82.83$ $43.40$ $69.90$ 18.80 $53.73$
TaskRes (CVPR’23) [31] $79.90$ $38.10$ $69.40$ $12.87$ $50.07$
MaPLe (CVPR’23) [15] $92.80$ $55.53$ $89.93$ $13.90$ $63.04$
PromptSRC (ICCV’23) [16] $92.55$ $55.17$ $91.40$ $14.83$ $63.49$
CLAP (CVPR’24) [25] $82.96$ $49.43$ $78.27$ $17.47$ $57.03$
TCP (CVPR’24) [29] $84.93$ $52.83$ $80.63$ $16.53$ $58.73$
@@
Table 2: Performance comparison (average accuracy (%) over 3 seeds) of the proposed MIST with the state-of-the-art approaches in the pCDFSL setting for $k=1,2,4,8,16$ shots from each class.
Inference: Once we learn the parameters of the distribution, during inference, we can now sample weights for the two text prompts as follows: $\theta_{t}^{1}\sim\mathcal{N}(\bar{\mu}_{0},\sigma_{0}^{*})$ and $\theta_{t}^{2}\sim\mathcal{N}(\mu_{1}^{*},\sigma_{1}^{*})$. For each class, we take the maximum logit among the two text prompts as the output prediction for that class.
@@ 5 Experimental Results
Datasets used: The proposed pCDFSL setting is source-free. For the target datasets, we consider the BSCDFSL [10] benchmark, consisting of four datasets, namely EuroSAT [11], ISIC [5], Plant Disease [22] and ChestX [27]. These datasets cover a varying spectrum of domain shifts, along with specialized classnames, encompassing satellite, agricultural and medical images. For training, we consider few samples ($k=1,2,4,8,16$) randomly selected from all the classes together and then evaluate the trained model on the full test set of all the datasets. The final accuracy is taken as the average over 3 different seeds.
@@
Implementation details: We employ CLIP ViT-B/16 as the backbone for our method similar to MaPLe [15]. The learnable context length of the text and vision inputs are set as $2$, and deep prompts are incorporated upto a depth of $9$. The model is trained using SGD optimizer for $150$ epochs with a learning rate of $0.0035$ and a batch size of $4$. All experiments are conducted on a NVIDIA RTX A5000 GPU.
5.1 Experiments in the proposed pCDFSL setting
@@
To validate the effectiveness of our approach, we compare our proposed MIST with several recent CLIP-based efficient transfer learning methods for varying number of shots. Specifically, we compare with 1) CoOp [32] and TCP [29] which employ prompt tuning on the text branch; 2) MaPLe [15] and PromptSRC [16] which utilizes a multimodal prompt tuning approach; 3) TaskRes [31] where task-specific adapters are tuned keeping the base text classifier fixed; 4) CLAP [25] uses a linear-probing approach and mainly addresses the absence of validation sets in FSL.
@@
For fair comparison, we run all the methods (using the official, publicly available codes) on the ViT-B/16 backbone and report the results in Table 2. We list some observations here: (i) Among the competing methods, MaPLe and PromptSRC achieves the highest performance on average, closely followed by TCP. Their improved performance suggests the effectiveness of multimodal prompt tuning in handling distribution shifts over text prompt tuning, which was also observed in [15]. This observation is further supported by our results; (ii) Although, CLAP is a recent approach, it mainly focuses on the validation problem of FSL. The reduced performance of CLAP highlights the limitations of linear probing, which does not utilize the text information for handling significant semantic and domain shifts; (iii) As the number of shots increases, the multimodal prompt tuning approaches like MaPLe, PromptSRC and MIST outperforms other methods by larger margins, suggesting that training more parameters is more effective for higher shots. (iv) Although all the methods show similar performance on the ChestX dataset, their overall accuracies are extremely low due to the large domain shift. However, text prompt tuning methods perform slightly better than the multimodal counterparts. This maybe because ChestX contains greyscale images, where additional prompt tuning in the vision branch degrades the performance. Overall, our proposed MIST framework outperforms the other methods significantly, giving consistent average gains of $3.19\%$, $2.01\%$, $1.79\%$, $3.19\%$, $1.60\%$ on $k=1,2,4,8,16$ shots respectively, over the best performing methods. The significant improvement for the 1-shot case highlights the effectiveness of our approach in mitigating overfitting in extremely low data scenarios.
@@ 5.2 Additional Analysis
Here we perform additional analysis and ablation studies to further validate our proposed framework. For the analysis, we compare with MaPLe, PromptSRC since they use multimodal prompts and are better suited for the pCDFSL task and TCP, since it is the state-of-the-art prompt tuning approach on CLIP.
1) Class-imbalanced learning: Here, we explore an even more challenging and realistic scenario, where the proposed pCDFSL setting may have varying training data from individual classes. The standard few-shot settings in literature assume an idealistic scenario where each class has exactly $k$ training samples, overlooking the effect of class imbalance. To create such a setting, we perform data sampling in a cyclic manner, e.g., we take $\{1,2,4,8,1,2,...\}$ from each class of the target dataset for training. The model is then evaluated on the entire test set. The results on two representative datasets, EuroSAT and ISIC in Table 3 shows that the proposed MIST outperforms the other methods even under class-imbalanced conditions, highlighting its effectiveness.
@@
MIST (Ours) $77.90$$\pm\textbf{2.63}$ 34.40 ±$\pm\textbf{3.56}$
@@
Table 4: The proposed MIST has much lower variance across three different random seeds compared to the other approaches.
@@
4) Ablation Study: Our proposed MIST framework models the two text classifiers using two distinct Gaussian sampling techniques as described earlier. Here we analyze the effectiveness of each of the proposed components in Table 5. Our base method is the existing multimodal prompt tuning framework, i.e., MaPLe. Introducing stochasticity to this single prompt (fixed mean) tuning approach by sampling from a learnable Gaussian distribution mitigates overfitting as described earlier, and improves the performance. Introducing the second learnable text prompt sampled from a fully learnable distribution without the $\mathcal{L}_{reg}$ term increases performance in EuroSAT, but reduces for ISIC. This can happen due to collapse of the two classifiers without the regularization term [1, 26]. Finally, adding the regularization term $\mathcal{L}_{reg}$ outperforms the baseline significantly for both the datasets.
@@
Table 5: Ablation study (1-shot): Each proposed component contributes to the impressive performance of the proposed framework.
Figure 5: Effect of prompt length (left) and number of text prompts per class (right). We report results for the EuroSAT (1-shot) case.
Number of text prompts & Prompt Length: MIST utilizes two prompts per class sampled from learnable Gaussian distributions. Here we study the effect of adding more text prompts to our approach. For this, we keep one prompt with a fixed mean, and all the others sampled from fully learnable distributions.
From Figure 5 (right), we observe that the performance starts decreasing after three prompts, suggesting overfitting from the increasing number of learnable parameters. Further, addition of more classifiers introduces increased computational overhead and longer training time. The effect of increasing the number of learnable prompt vectors is illustrated in Figure 5 (left). Here also, the accuracy decreases after a point, indicating overfitting in the few-shot setting. We have used 2 learnable prompts for all the experiments.
@@
Figure 6: Average accuracies for the pCDFSL and CLIP-based FSL benchmark datasets for the 4-shot setting. MIST outperforms other methods in the pCDFSL setting, and is second to TCP in FSL. Overall, MIST performs better on all the datasets combined.
@@
5) Future Directions: Existing CDFSL methods in literature typically do not evaluate on natural image datasets and FSL methods do not test on significant domain shifts. However, it is important for the research community to develop more generalized frameworks which work across diverse semantics and distribution shifts seamlessly. Although MIST is designed to address the pCDFSL task, we study its effect on natural image datasets with simpler domain or semantic shifts. For this, we consider the CLIP-based FSL datasets like DTD, Caltech101, Oxford Pets, Flowers, Aircraft, Food101. From Figure 6, we observe that the proposed MIST is only second to the state-of-the-art TCP [29] on average for these datasets and outperforms all the other approaches. It is worth mentioning that unlike MIST, 1) TCP and PromptSRC utilizes detailed class descriptions to incorporate prior knowledge, which requires handcrafting well-designed prompts; 2) they also require the zero-shot CLIP model, resulting in increased memory consumption. Overall, the proposed MIST outperforms all the other methods on average when considered across all datasets encompassing both natural and domain-shifted images and classnames, as observed in Figure 6.

### v2

@@ 4.2 MIST: Multiple Stochastic Prompt Tuning
Since CLIP is pretrained on web-scale data encompassing mainly natural images clip , the image encoder struggles to learn robust classwise features when faced with extreme domain and semantic shifts in the target dataset. In addition, intraclass diversity and interclass similarity in the images due to presence of many classes, results in disjoint clusters in visual features from the same class. Consider an illustrative example of the EuroSAT dataset containing satellite images of various terrains. Here, distinct classes like “Highway” and “River” can have similar visual representatives in the few-shot training data, at the same time featuring diverse visuals from the same classes as shown in Fig. 2. Existing prompt tuning approaches represent each class with single learnable prompts, implicitly assuming that classwise visual features form single clusters. However, such a strategy is insufficient to represent the disjoint visual clusters, which may result from such challenging settings. To illustrate this, we consider two representative prompt-tuning methods, MaPLe maple and TCP tcp and show their t-SNE visualizations in Fig. 2. We observe that the unimodal assumption is violated and the image embeddings of each class form multiple modes in the representation space.
To address this, inspired from imp ; mixtfsl , we introduce a multiple prompt learning approach, where we represent each class with multiple prompt vectors. However, incorporating too many learnable prompts for each class is not desirable since: (i) Each class contains few training samples, and many text embeddings per class could lead to overfitting on individual datapoints, resulting in loss of class level representations, and (ii) it introduces additional learnable parameters and thus more computational overhead. As a balanced approach, we represent each class with two prompt vectors. Formally, let the embeddings corresponding to the two prompts for a particular class $CLS_{k}$ be denoted as $\tilde{z}_{t}^{i}$, where, $i$ = 1, 2. Here, $\tilde{z}_{t}^{i}=\mathcal{F}_{t}(\tilde{X}_{t}^{i})$, where, $\tilde{X}_{t}^{i}=\{t_{SOS},\theta_{t}^{i},t_{1},...,t_{CLS_{k}},t_{EOS}\}$ is the $i^{th}$ text prompt for the class $CLS_{k}$. $\theta_{t}^{i}=\{\theta_{t1}^{i},\theta_{t2}^{i},...,\theta_{tm}^{i}\}$ denotes the $i^{th}$ set of learnable prompt vectors for that particular class.
@@
CoOp (IJCV’22) $51.87$ $22.77$ $24.73$ 22.83 $30.55$
TaskRes (CVPR’23) $64.67$ $19.70$ $36.57$ $10.97$ $32.98$
MaPLe (CVPR’23) $73.30$ $27.50$ $51.53$ $14.60$ $41.73$
PromptSRC (ICCV’23) $73.23$ $21.97$ 55.03 $14.37$ $41.15$
CLAP (CVPR’24) $61.46$ $26.61$ $47.22$ $15.94$ $37.81$
TCP (CVPR’24) $64.30$ $27.80$ $49.37$ $14.93$ $39.10$
@@
2-shot
CoOp (IJCV’22) $66.00$ $21.87$ $37.97$ $14.43$ $35.07$
TaskRes (CVPR’23) $68.83$ $23.13$ $39.27$ $10.83$ $35.52$
MaPLe (CVPR’23) $78.07$ $31.90$ $67.17$ $16.27$ $48.35$
PromptSRC (ICCV’23) $79.53$ $29.47$ $68.07$ $12.70$ $47.44$
CLAP (CVPR’24) $70.63$ $34.79$ $60.13$ 16.43 $45.50$
TCP (CVPR’24) $70.37$ 36.87 $62.63$ $15.63$ $46.38$
@@
4-shot
CoOp (IJCV’22) $66.53$ $25.00$ $42.67$ $17.93$ $38.03$
TaskRes (CVPR’23) $72.40$ $21.40$ $39.35$ $10.27$ $35.86$
MaPLe (CVPR’23) $84.03$ $37.17$ $77.07$ 19.73 $54.50$
PromptSRC (ICCV’23) $85.23$ $37.63$ $78.70$ $15.17$ $54.18$
CLAP (CVPR’24) $76.43$ $34.37$ $65.11$ $18.98$ $48.72$
TCP (CVPR’24) $76.77$ $37.37$ $67.97$ $17.07$ $49.80$
@@
8-shot
CoOp (IJCV’22) $76.53$ $38.27$ $60.50$ $14.60$ $47.48$
TaskRes (CVPR’23) $74.63$ $34.30$ $57.77$ $12.57$ $44.82$
MaPLe (CVPR’23) $86.80$ $46.53$ $84.47$ $14.17$ $57.99$
PromptSRC (ICCV’23) $88.37$ $42.47$ $86.80$ $14.93$ $58.14$
CLAP (CVPR’24) $76.85$ $42.81$ $74.62$ $14.97$ $52.31$
TCP (CVPR’24) $79.03$ $46.57$ $76.33$ $14.97$ $54.23$
@@
16-shot
CoOp (IJCV’22) $82.83$ $43.40$ $69.90$ 18.80 $53.73$
TaskRes (CVPR’23) $79.90$ $38.10$ $69.40$ $12.87$ $50.07$
MaPLe (CVPR’23) $92.80$ $55.53$ $89.93$ $13.90$ $63.04$
PromptSRC (ICCV’23) $92.55$ $55.17$ $91.40$ $14.83$ $63.49$
CLAP (CVPR’24) $82.96$ $49.43$ $78.27$ $17.47$ $57.03$
TCP (CVPR’24) $84.93$ $52.83$ $80.63$ $16.53$ $58.73$
@@
Table 2: Performance comparison (average accuracy (%) over 3 seeds) of the proposed MIST with the state-of-the-art approaches for $k=1,2,4,8,16$ shots from each class.
Inference: After learning the parameters of the distribution, during inference, we can sample weights for the two text prompts as follows: $\theta_{t}^{1}\sim\mathcal{N}(\bar{\mu}_{0},\sigma_{0}^{*})$ and $\theta_{t}^{2}\sim\mathcal{N}(\mu_{1}^{*},\sigma_{1}^{*})$. For each class, we take the maximum logit among the two text prompts as the output prediction for that class.
@@ 5 Experimental Results
Datasets used: For the target datasets, we consider the BSCDFSL bscdfsl benchmark, which is collected from real-world settings, and consists of four datasets, namely EuroSAT eurosat , ISIC isic , Plant Disease plant_disease and ChestX chestx . These datasets cover a varying spectrum of domain shifts, along with specialized classnames, encompassing satellite, agricultural and medical images. For training, we consider few samples ($k=1,2,4,8,16$) randomly selected from all the classes together and then evaluate the trained model on the full test set of all the datasets. The final accuracy is taken as the average over 3 different seeds.
@@
Implementation details: We employ CLIP ViT-B/16 as the backbone similar to MaPLe maple . The learnable context length of the text and vision inputs are set as $2$, and deep prompts are incorporated upto a depth of $9$. The model is trained using SGD optimizer for $150$ epochs with a learning rate of $0.0035$ and a batch size of $4$. All experiments are conducted on a NVIDIA RTX A5000 GPU.
5.1 Comparison with state-of-the-art methods
@@
To validate the effectiveness of our approach, we compare our proposed MIST with several recent CLIP-based efficient transfer learning methods for varying number of shots. Specifically, we compare with 1) CoOp coop and TCP tcp , which employ prompt tuning on the text branch; 2) MaPLe maple and PromptSRC promptsrc which utilize a multimodal prompt tuning approach; 3) TaskRes taskres where task-specific adapters are tuned keeping the base text classifier fixed; 4) CLAP clap uses a linear-probing approach and mainly addresses the absence of validation sets in FSL.
For fair comparison, we run all the methods (using the official, publicly available codes) on the ViT-B/16 backbone and report the results in Table 2. We list some observations below:
(i) Among the competing methods, MaPLe and PromptSRC achieves the highest performance on average, closely followed by TCP. Their improved performance suggests the effectiveness of multimodal prompt tuning in handling distribution shifts over text prompt tuning, which was also observed in maple . This observation is further supported by our results;
(ii) Although, CLAP is a recent approach, it mainly focuses on the validation problem of FSL. The reduced performance of CLAP highlights the limitations of linear probing, which does not utilize the text information for handling significant semantic and domain shifts;
(iii) As the number of shots increases, the multimodal prompt tuning approaches like MaPLe, PromptSRC and MIST outperforms other methods by larger margins, suggesting that training more parameters is more effective for higher shots.
@@
(iv) Although all the methods show similar performance on the ChestX dataset, their overall accuracies are extremely low due to the large domain shift. However, text prompt tuning methods perform slightly better than the multimodal counterparts. This maybe because ChestX contains greyscale images, where additional prompt tuning in the vision branch degrades the performance. Overall, our proposed MIST framework outperforms the other methods significantly, giving consistent average gains of $3.19\%$, $2.01\%$, $1.79\%$, $3.19\%$, $1.60\%$ on $k=1,2,4,8,16$ shots respectively, over the best performing methods. The significant improvement for the 1-shot case highlights the effectiveness of our approach in mitigating overfitting in extremely low data scenarios.
@@ 5.2 Additional Analysis
Here we perform additional analysis and ablation studies to further validate our proposed framework. For the analysis, we compare with MaPLe, PromptSRC since they use multimodal prompts and are better suited for this task and TCP, since it is the state-of-the-art prompt tuning approach on CLIP.
1) Class-imbalanced learning: Here, we explore an even more challenging scenario, where the number of labeled examples may vary across classes, reflecting real-world datasets. The standard few-shot settings in literature assume an idealistic scenario where each class has exactly $k$ training samples, overlooking the effect of class imbalance. To create such a setting, we perform data sampling in a cyclic manner, e.g., we take $\{1,2,4,8,1,2,$...\} from each class of the target dataset for training. The model is then evaluated on the entire test set. The results on two representative datasets, EuroSAT and ISIC in Table 3 shows that the proposed MIST outperforms the other methods even under class-imbalanced conditions, highlighting its effectiveness.
@@
MIST (Ours) $77.90$ $\pm\textbf{2.63}$ 34.40 $\pm\textbf{3.56}$
@@
Table 4: MIST exhibits significantly lower variance across three different random seeds compared to other approaches.
@@
4) Ablation Study: Our proposed MIST framework models the two text classifiers using two distinct Gaussian sampling techniques as described earlier. Here we analyze the effectiveness of each of the proposed components in Table 5. Our base method is the multimodal prompt tuning framework, MaPLe. Introducing stochasticity to this single prompt (fixed mean) tuning approach by sampling from a learnable Gaussian distribution mitigates overfitting as described earlier, and improves performance. Introducing the second learnable prompt sampled from a fully learnable distribution without the $\mathcal{L}_{reg}$ term increases performance in EuroSAT, but reduces for ISIC. This can happen due to collapse of the two classifiers without the regularization term mixtfsl ; deepfake . Finally, adding the regularization term $\mathcal{L}_{reg}$ outperforms the baseline significantly for both the datasets.
@@
Table 5: Ablation study (1-shot): All the proposed components collectively enhance the overall performance.
Figure 5: Effect of prompt length (left) and number of text prompts per class (right) for EuroSAT data (1-shot).
5) Number of text prompts & Prompt Length: MIST utilizes two prompts per class sampled from learnable Gaussian distributions. Here we study the effect of adding more text prompts. For this, we keep one prompt with a fixed mean, and the others sampled from fully learnable distributions. From Figure 5 (right), we observe that the performance starts decreasing after three prompts, suggesting overfitting from the increasing number of learnable parameters. Further, addition of more classifiers introduces increased computational overhead and longer training time. The effect of increasing the number of learnable prompt vectors is illustrated in Figure 5 (left). Here also, the accuracy decreases after a point, indicating overfitting in the few-shot setting. We used 2 learnable prompts for all the experiments.
@@
Limitations. While MIST outperforms state-of-the-art methods across all datasets, its performance slightly drops on the grayscale ChestX dataset, likely because of the additional visual prompts. In such cases, methods relying solely on textual prompts may prove more effective.